# Improved Streaming Algorithm for Fair $k$-Center Clustering

### v1

@@ Abstract
Many real-world applications pose challenges in incorporating fairness constraints into the $k$-center clustering problem, where the dataset consists of $m$ demographic groups, each with a specified upper bound on the number of centers to ensure fairness. Focusing on big data scenarios, this paper addresses the problem in a streaming setting, where data points arrive one by one sequentially in a continuous stream. Leveraging a structure called the $\lambda$-independent center set, we propose a one-pass streaming algorithm that first computes a reserved set of points during the streaming process. Then, for the post-streaming process, we propose an approach for selecting centers from the reserved point set by analyzing all three possible cases, transforming the most complicated one into a specially constrained vertex cover problem in an auxiliary graph. Our algorithm achieves a tight approximation ratio of 5 while consuming $O(k\log n)$ memory. It can also be readily adapted to solve the offline fair $k$-center problem, achieving a 3-approximation ratio that matches the current state of the art. Furthermore, we extend our approach to a semi-structured data stream, where data points from each group arrive in batches. In this setting, we present a 3-approximation algorithm for $m=2$ and a 4-approximation algorithm for general $m$. Lastly, we conduct extensive experiments to evaluate the performance of our approaches, demonstrating that they outperform existing baselines in both clustering cost and runtime efficiency.
@@ 1 Introduction
Fair $k$-center clustering is a popular problem in various fields including data summarization (Kleindessner, Awasthi, and Morgenstern 2019; Angelidakis et al. 2022) and machine learning (Chierichetti et al. 2017; Jones, Nguyen, and Nguyen 2020). The problems have been widely studied, with many definitions of fairness proposed and corresponding approximation algorithms discussed, such as group fairness (Chierichetti et al. 2017; Wu et al. 2024; Backurs et al. 2019), data summarization fairness (Kleindessner, Awasthi, and Morgenstern 2019; Jones, Nguyen, and Nguyen 2020; Wu et al. 2024; Lin, Guo, and Jia 2024), colorful fairness (Bandyapadhyay et al. 2019; Anegg, Vargas Koch, and Zenklusen 2022; Jia, Sheth, and Svensson 2022) and so on. In this paper, we focus on the data summarization fairness $k$-center problem. We consider a dataset $S$ of size $n$ divided into $m$ disjoint groups, as $S=\{S_{1}\cup\ldots\cup S_{m}\}$. Our goal is to select $k$ clusters to minimize the maximum distance from each point to its nearest center, while the number of centers chosen from each group $S_{l}$ is bound by $k_{l}$. This formulation can avoid the biased for some sensitive features by controlling the number of objects from each category in the output. For instance, it can dictate the number of movies from each genre shown to a user in a recommendation system or limit the number of old messages included in a summary of a user’s feed (Mahabadi and Trajanovski 2024).
However, a traditional challenge in clustering is the need to process large-scale datasets across numerous applications. In such scenarios, storing the entire input in memory becomes impractical, giving rise to the streaming model. In this model, data points arrive sequentially in a stream, and only a limited portion can be retained in memory due to space constraints. The streaming algorithm must decide, upon the arrival of each data point, whether to store it or discard it. Notably, designing effective streaming algorithms is more challenging than developing offline algorithms, as decisions must be made based on partial information rather than having access to the entire dataset. Motivated by this challenge, we aim to propose the approximation algorithms for the fair $k$-center problem under the streaming setting.
@@ 1.1 Related Work
To the best of our knowledge, the fair $k$-center clustering problem was first formally addressed in the context of data summarization by Kleindessner, Awasthi, and Morgenstern (2019). They proposed a $5$-approximation algorithm for the case of two groups and a $(3\times 2^{m-1}-1)$-approximation for the general case with $m$ groups. After them, the improvement was made by Jones, Nguyen, and Nguyen (2020), who reduced the approximation ratio to $3$ for arbitrary $m$ groups. Prior to these two relevant works, the matroid center problem which includes a generalization of the $k$-center problem that enforces a matroid constraint on the set of centers rather than a simple cardinality constraint, was studied by Chen et al. (2016). They developed a $3$-approximation algorithm, although with considerably higher time complexity. More recently, Chen et al. (2024) introduced the fair $k$-supplier problem, which involves selecting $k$ facilities from a dataset partitioned into $m$ disjoint groups, subject to group-wise upper bounds on the number of facilities selected. They presented a $5$-approximation algorithm for this problem by maximum matching techniques.
@@
For large-scale datasets in real-time streaming scenarios, streaming $k$-center clustering was first studied by Matthew McCutchen and Khuller (2008), which provided a streaming $(4+\epsilon)$-approximation algorithm with outliers, where up to $z$ input points can be dropped and $O(kz/\epsilon)$ memory was consumed. Then, a deterministic one-pass streaming algorithm was developed for the same problem by Ceccarello, Pietracaprina, and Pucci (2019), achieving an improved approximation ratio $(3+\epsilon)$ with $O\big((k+z)(96/\epsilon)^{d}\big)$ memory with $d$-dimension. For the matroid center problem, a $(17+\epsilon)$-approximation one-pass algorithm with a running time $O_{\epsilon}\big((nk+k^{3.5})+k^{2}\log(\Lambda)\big)$ has been developed by Kale (2019), where $k$ is the rank of the matroid, $\Lambda$ is the aspect ratio of the metric, and $\epsilon$ terms are hidden by the $O_{\epsilon}$ notation. For the fair $k$-center problem, Chiplunkar, Kale, and Ramamoorthy (2020) provided a distributed algorithm achieving $(17+\epsilon)$ approximation ratio with $O(kn/l+mk^{2}l)$ running time for $l$ processors. They also developed a two-pass streaming algorithm with approximation ratio $3$ for the fair $k$-center problem. Recently, the fair range $k$-center was addressed by Nguyen, Nguyen, and Jones (2022) with an achieved approximation ratio $13$. In addition, a one-pass streaming algorithm for fair $k$-center clustering proposed by Lin, Guo, and Jia (2024) achieved the state-of-the-art approximation ratio of 7. In this paper, we improve the approximation ratio to 5 by utilizing the independent centers and modifying the post-streaming algorithms.
@@ 1.2 Our Contributions
The main contribution of this paper is summarized as follows:
  • Devise a one-pass streaming algorithm for fair $k$-center, achieving an improved approximation ratio of $5$ while consuming $O(k\log n)$ memory. This significantly improves the previous SOTA ratio $7$ due to Lin, Guo, and Jia (2024).
  • For semi-structured data streams where data points of each group $S_{l}$ are streamed as a batch, we achieve ratios 3 and 4 for $m=2$ and general $m$, respectively.
  • Construct extensive experiments on real-world datasets to demonstrate the practical performance gains of our algorithms in both clustering accuracy and runtime efficiency.
Notably, we also provide an example to show that the achieved ratio 5 is tight under the $o(n)$ memory constraint. Moreover, the streaming algorithm with ratio $5$ can be slightly tuned to achieve the SOTA ratio 3 for offline fair $k$-center previously due to Jones, Nguyen, and Nguyen (2020).
@@ 2 Preliminary and Algorithmic Framework
Let $S$ be a finite set of data points with size $n$ distributed in the metric space where $d:S\times S\rightarrow\mathbb{R}_{\geq 0}$ is the distance on $S$ that satisfies the triangle inequality. For a given parameter $k\in\mathbb{N}$, the traditional $k$-center clustering problem is to select a set of $k$ points $C\subseteq S$ to serve all points in $S$ such that $\max_{s\in S}d\left(s,C\right)$ is minimized, where $d\left(s,C\right)=\min_{c\in C}d\left(s,c\right)$ is the distance between a point $s$ and the center set $C$. Let the data set be divided into $m$ disjoint groups $S=\{S_{1}\cup\ldots\cup S_{m}\}$, where there exists a fairness constraint that the number of chosen centers from group $S_{l}$ is constrained by an upper bound $k_{l}$, where $k_{l}$ is the fair constraint for each group $l$ for $l\in[m]$ and $\sum^{m}_{l=1}k_{l}=k$, where we use $m]=\{1,2,\dots,m\}$ for briefness. Then, the fair $k$-center clustering problem is to find a center set $C$ satisfying the formulation as follows:
$\min_{C\subseteq S}$ $\max_{s\in S}d\left(s,C\right)$
$s.t.$ $\mid C\cap S_{l}\mid\leq k_{l},\,\forall l$
$\mid C\cap S\mid=\sum_{l=1}^{m}\mid C\cap S_{l}\mid\leq k.$
@@ 2.1 Algorithmic Framework
In this subsection, we introduce the general framework of our algorithms. Throughout this paper, we assume that the optimal radius of the fair $k$-center problem is known as $r^{*}$ (i.e., $r^{*}=\max_{s\in S}d\left(s,C^{*}\right)$ for the optimum center set $C^{*}$). Although the exact value of $r^{*}$ is actually unknown, we can employ the previous approach for finding a suitable replacement of $r^{*}$ due to Guo et al. (2025). In general, all of our algorithms mainly proceed in two stages.
  • Streaming stage. Select a set of representative data points along the stream according to the optimum radius $r^{*}$, where the selected set might be with a size larger than $k$;
  • Post-streaming Stage. Compute the set of actually centers from the set of streamed and stored representative data points according to $r^{*}$.
@@ 2.2 The streaming stage
For the first stage, we employ the $\lambda$-independent center set as defined below, but simply tuned the streaming stage for different ratios:
@@
($\lambda$-independent center set) $\Gamma\subseteq S$ is a $\lambda$-independent center set of $S$, if and only if it satisfies the following two conditions:
@@
We say a $\lambda$-independent center set $\Gamma\subseteq S$ is minimal iff removal of any point from $\Gamma$ makes $\Gamma$ no longer a $\lambda$-independent center set. Assume $C^{*}$ is the center set of an optimal solution and recalled that $r^{*}$ is the optimal radius. Then we have:
@@
We need only to show the case for $\lambda=2r^{*}$. Suppose $|\Gamma|>k$. Then, by the Pigeonhole principle, there must exist at least a pair of points $i$, $j\in\Gamma$, which are covered by the same center in the optimal solution, say $c^{*}$ in $C^{*}$. Then, both $d(i,c^{*})\leq r^{*}$ and $d(c^{*},j)\leq r^{*}$ hold. By the triangle inequality, we then have $d(i,j)\leq d(i,c^{*})+d(c^{*},j)\leq r^{*}+r^{*}=2r^{*}$. On the other hand, we have $d(i,j)>2r^{*}$ according to Cond. (1) in Def. 1, resulting in a contradiction and hence completing the proof. ∎
3 Fair k-center Clustering in Data Streams
In this section, we present our two-stage streaming algorithm for the case $m=2$, and then extend to general $m$.
3.1 Overview of the algorithm
For Stage 1, we need only to simply construct a minimal $\lambda$-independent center set $\Gamma_{l}$ for each $l\in\{1,2\}$ regarding $\lambda=2r^{*}$ along the stream:
  • Upon each arriving point $i$, if $i\in S_{l}$ and $d\left(i,\Gamma_{l}\right)>\lambda$ both hold, grow $\Gamma_{l}$ (initially empty) by adding $i$.
Next, we present details of Stage 2 with a ratio proof based on analyzing the sizes of $\Gamma_{1}$ and $\Gamma_{2}$. There are exactly three possibilities regarding the sizes of $\Gamma_{1}$ and $\Gamma_{2}$:
@@ 3.1 Overview of the algorithm
For Case (1), we can directly use $C=\Gamma_{1}\cup\Gamma_{2}$ as the desired center set, because we have: (1) $|C|=|\Gamma_{1}\cup\Gamma_{2}|\leq k_{1}+k_{2}=k$; (2) for any point $s\in S$, $d(s,C)\leq 2r^{*}$ holds due to the definition of $\lambda$-independent center set and $\lambda=2r^{*}$.
Then, we will give a method to solve Case (2), and show that Case (3) can be reduced to Case (2).
3.2 Procession of Case (2)
Without loss of generality, we assume that $|\Gamma_{l}|>k_{l}$ and $|\Gamma_{3-l}|\leq k_{3-l}$ for $l\in\{1,2\}$. Our algorithm constructs a new set $\Gamma_{l}^{\prime}=\{i|i\in\Gamma_{l},d\left(i,\Gamma_{3-l}\right)>3r^{*}\}$, and simple uses $C=\Gamma_{3-l}\cup\Gamma_{l}^{\prime}$ as the center set. The correctness of the such $C$ can be derived from the following theorem:

### v2

@@ Abstract
Many real-world applications call for incorporating fairness constraints into the $k$-center clustering problem, where the dataset is partitioned into $m$ demographic groups, each with a specified upper bound on the number of centers to ensure fairness. Focusing on big data scenarios, this paper addresses the problem in a streaming setting, where data points arrive sequentially in a continuous stream. Leveraging a structure called the $\lambda$-independent center set, we propose a one-pass streaming algorithm that first computes a reserved set of points during the streaming process. In the post-streaming process, we then select centers from the reserved point set by analyzing three possible cases and transforming the most complex one into a specially constrained vertex-cover problem on an auxiliary graph. Our algorithm achieves an approximation ratio of $5+\epsilon$ and memory complexity $O(k\log\alpha)$, where $\alpha$ is the aspect ratio and $\epsilon>0$ is any small constant. Furthermore, we extend our approach to semi-structured data streams, where data points arrive in groups. In this setting, we present a $(3+\epsilon)$-approximation algorithm for $m=2$, which can be readily adapted to solve the offline fair $k$-center problem, achieving an approximation ratio of 3 that matches the current state of the art. Lastly, we conduct extensive experiments to evaluate the performance of our approaches, demonstrating that they outperform existing baselines in both clustering cost and runtime efficiency.
Codehttps://github.com/ChaoqiJia/StreamFKC
@@ 1 Introduction
Fair $k$-center clustering is a popular problem in various fields, including data summarization (Kleindessner et al. 2019; Angelidakis et al. 2022) and machine learning (Chierichetti et al. 2017; Jones et al. 2020). The problem has been widely studied, with many definitions of fairness proposed and corresponding approximation algorithms discussed, such as group fairness (Chierichetti et al. 2017; Wu et al. 2024; Backurs et al. 2019), data summarization fairness (Kleindessner et al. 2019; Ceccarello et al. 2024; Wu et al. 2024; Lin et al. 2024), colorful fairness (Bandyapadhyay et al. 2019; Anegg et al. 2022; Jia et al. 2022) and so on. In this paper, we focus on the fair $k$-center problem for data summarization. We consider a dataset $S$ of size $n$ divided into $m$ disjoint groups, denoted as $S=S_{1}\cup\ldots\cup S_{m}$. Our goal is to select $k$ centers to minimize the maximum distance from each point to its nearest center, while the number of centers chosen from each group $S_{l}$ is bounded by $k_{l}$. This formulation helps avoid bias toward some sensitive features by controlling the number of objects from each category in the output. For example, it can dictate the number of movies of each genre shown to a user in a recommendation system or limit the number of old messages included in a summary of a user’s feed (Mahabadi and Trajanovski 2024).
A traditional challenge in clustering is the need to process large-scale datasets across numerous applications. In such scenarios, storing the entire input in memory becomes impractical, giving rise to the streaming model. In this model, data points arrive sequentially in a stream, and only a limited portion can be retained in memory due to space constraints. The streaming algorithm must decide, upon the arrival of each data point, whether to store it or discard it. Notably, designing effective streaming algorithms is more challenging than developing offline algorithms, as decisions must be made based on partial information rather than having access to the entire dataset. Motivated by this challenge, we aim to propose approximation algorithms for the fair $k$-center problem under the streaming setting.
@@ 1.1 Related Work
Fairness is an important concept for many research fields such as data summarization (Kleindessner et al. 2019), private security (Zhang et al. 2024b), and machine learning (Zhang et al. 2024a, 2025). In this paper, we briefly review the literature on the $k$-center problem, focusing on studies about data summarization fairness and streaming model settings.
@@
The fair $k$-center clustering was formally addressed in the context of data summarization by Kleindessner et al. (2019). They proposed a $5$-approximation algorithm for the case of two groups and a $(3\times 2^{m-1}-1)$-approximation for the general case with $m$ groups. Subsequently, Jones et al. (2020) reduced the approximation ratio to $3$ for arbitrary $m$ groups. Before these two relevant works, the matroid center problem, a generalization of the $k$-center problem that enforces a matroid constraint on the center set rather than a simple cardinality constraint, was studied by Chen et al. (2016). They developed a $3$-approximation algorithm, although with considerably higher time complexity. More recently, Chen et al. (2024) introduced the fair $k$-supplier problem, which involves selecting $k$ facilities from a dataset partitioned into $m$ disjoint groups, subject to group-wise upper bounds on the number of facilities selected. They presented a $5$-approximation algorithm for this problem by maximum matching techniques.
@@
For large-scale datasets in real-time streaming scenarios, streaming $k$-center clustering was studied by Matthew McCutchen and Khuller (2008), who provided a streaming $(4+\epsilon)$-approximation algorithm with outliers, where up to $z$ input points can be dropped and $O(kz/\epsilon)$ memory is used. Then, a deterministic one-pass streaming algorithm was developed for the same problem by Ceccarello et al. (2019), achieving an improved approximation ratio $(3+\epsilon)$ with $O\big((k+z)(96/\epsilon)^{d}\big)$ memory in $d$ dimensions. For the matroid center problem, a $(17+\epsilon)$-approximation one-pass algorithm with a running time $O_{\epsilon}\big((nk+k^{3.5})+k^{2}\log(\Lambda)\big)$ has been developed by Kale (2019), where $k$ is the rank of the matroid, $\Lambda$ is the aspect ratio of the metric, and $\epsilon$ terms are hidden by the $O_{\epsilon}$ notation. For the fair $k$-center problem, Chiplunkar et al. (2020) provided a distributed algorithm achieving an approximation ratio of $(17+\epsilon)$ and with $O(kn/l+mk^{2}l)$ running time for $l$ processors. They also developed a two-pass streaming algorithm with an approximation ratio $3$ for the fair $k$-center problem. Recently, the fair range $k$-center was addressed by Nguyen et al. (2022) with an approximation ratio of $13$ and Ceccarello et al. (2024) employed a coreset-based method for the colorful fair $k$-center problem in doubling metrics, achieving a $(3+\epsilon)$-approximation ratio for sequential, streaming and MapReduce/MPC algorithms. In addition, a one-pass streaming algorithm for representative fair $k$-center clustering by Lin et al. (2024) achieved the state-of-the-art approximation ratio of $(7+\epsilon)$. This paper improves the approximation ratio to $(5+\epsilon)$ under a streaming setting by utilizing $\lambda$-independent center sets and modifying the post-streaming algorithms.
@@ 1.2 Our Contributions
The main contributions can be summarized as follows:
  • Devise a one-pass streaming algorithm for fair $k$-center, achieving an improved approximation ratio of $5+\epsilon$ while consuming $O(k\log\alpha)$ memory, where $\alpha$ is the aspect ratio and $\epsilon>0$ is any small constant. This significantly improves the previous SOTA ratio $7+\epsilon$ due to Lin et al. (2024).
  • For semi-structured data streams where data points of each group $S_{l}$ are streamed as a batch, we propose a streaming algorithm achieving an approximation ratio of $3+\epsilon$ for $m=2$.
  • Conduct extensive experiments on real-world datasets to demonstrate the practical performance gains of our algorithms in clustering accuracy and runtime efficiency.
Notably, we also provide an example demonstrating that our ratio analysis is tight for the algorithm. Moreover, the streaming algorithm with ratio $3+\epsilon$ can be adapted to solve the offline fair $k$-center, achieving a ratio of $3$, which matches the state-of-the-art ratio 3 for the offline setting due to Jones et al. (2020).
@@ 2 Preliminary and Algorithmic Framework
Let $S$ be a finite set of data points with size $n$ distributed in a metric space, where $d:S\times S\rightarrow\mathbb{R}_{\geq 0}$ is the distance on $S$ that satisfies the triangle inequality. For a given parameter $k\in\mathbb{N}$, the traditional $k$-center clustering problem is to select a set of $k$ points $C\subseteq S$ to serve all points in $S$ such that $\max_{s\in S}d\left(s,C\right)$ is minimized, where $d\left(s,C\right)=\min_{c\in C}d\left(s,c\right)$ is the distance between a point $s$ and the center set $C$. Let the data set be divided into $m$ disjoint groups $S=S_{1}\cup\ldots\cup S_{m}$. There exists a fairness constraint that the number of chosen centers from group $S_{l}$ is bounded by $k_{l}$, where $k_{l}$ is the upper bound for group $l$ for all $l\in[m]$ and $[m]$ denotes $\{1,2,\dots,m\}$ for brevity. Moreover, $\sum^{m}_{l=1}k_{l}=k$. Then, the fair $k$-center clustering problem is to find a center set $C$ satisfying the formulation as follows:
$\min_{C\subseteq S}\quad$ $\max_{s\in S}\;d(s,C)$
s.t. $\left|C\cap S_{l}\right|\leq k_{l},\quad\forall l,$
$\left|C\right|\leq k.$
@@ 2.1 Algorithmic Framework
Next, we introduce the general framework of our algorithms. Throughout this paper, we assume that the optimal radius of the fair $k$-center problem is known as $r^{*}$ (i.e., $r^{*}=\max_{s\in S}d\left(s,C^{*}\right)$ for the optimum center set $C^{*}$). Although the exact value of $r^{*}$ is actually unknown, we can employ the previous approach for finding suitable replacement of $r^{*}$ in the offline setting due to Guo et al. (2025) by compromising the ratio with a multiplicative factor of $1+\epsilon$. In general, all of our algorithms mainly proceed in two stages.
  • Streaming stage. Select a set of representative data points along the stream according to the optimum radius $r^{*}$, where the union of the selected point sets might have size larger than $k$;
  • Post-streaming stage. Compute the set of actual centers from the set of streamed and stored representative data points according to $r^{*}$.
@@ 2.2 The Streaming Stage
For the streaming stage, we employ the $\lambda$-independent center set as defined below:
@@
($\lambda$-independent center set) $\Gamma\subseteq S$ is a $\lambda$-independent center set of $S$ if and only if it satisfies the following two conditions:
@@
We say a $\lambda$-independent center set $\Gamma\subseteq S$ is minimal iff removal of any point from $\Gamma$ makes $\Gamma$ no longer a $\lambda$-independent center set. Assume $C^{*}$ is the center set of an optimal solution and recall that $r^{*}$ is the optimal radius. Then we have:
@@
We only need to show the case for $\lambda=2r^{*}$. Suppose $|\Gamma|>k$. Then, by the Pigeonhole principle, there must exist at least a pair of points $i$, $j\in\Gamma$, which are covered by the same center in the optimal solution, say $c^{*}$ in $C^{*}$. Then, both $d(i,c^{*})\leq r^{*}$ and $d(c^{*},j)\leq r^{*}$ hold. By the triangle inequality, we then have $d(i,j)\leq d(i,c^{*})+d(c^{*},j)\leq r^{*}+r^{*}=2r^{*}$. On the other hand, we have $d(i,j)>2r^{*}$ according to Cond. (1) in Def. 1, resulting in a contradiction and hence completing the proof. ∎
So for the streaming stage, we only need to construct a minimal $\lambda$-independent center set $\Gamma_{l}$ for each $l\in\{1,2\}$ regarding $\lambda=2r^{*}$ along the stream: Upon each arriving point $i$ do if $i\in S_{l}$ and $d\left(i,\Gamma_{l}\right)>\lambda$ both hold then Grow $\Gamma_{l}$ (initially empty) by adding $i$.
3 Post-streaming Stage of Fair $k$-Center
Observing that the size of $\Gamma_{1}\cup\Gamma_{2}$ obtained in the streaming stage may exceed $k_{l}$ for $\lambda=2r^{*}$, we devise a method to construct the actual center set $C$ by selecting points from the union of $\Gamma_{l}$ for $m=2$ based on dealing with the cases whether $\mid\Gamma_{l}\mid\leq k_{l}$ holds or not. We extend the method to general $m$ in the full version.
3.1 Overview of Post-streaming Stage
Our algorithm for this stage is grounded in analyzing the sizes of $\Gamma_{1}$ and $\Gamma_{2}$. There are exactly three possible cases regarding the sizes of $\Gamma_{1}$ and $\Gamma_{2}$:
@@ 3.1 Overview of Post-streaming Stage
For Case (1), we can directly use $C=\Gamma_{1}\cup\Gamma_{2}$ as the desired center set, because: (1) $|C|=|\Gamma_{1}\cup\Gamma_{2}|\leq k_{1}+k_{2}=k$; (2) for any point $s\in S$, $d(s,C)\leq 2r^{*}$ holds due to the definition of $\lambda$-independent center set and $\lambda=2r^{*}$.
Then, we give a method to solve Case (2) and show that Case (3) can be reduced to Case (2).
3.2 Processing of Case (2)
Without loss of generality, we assume that $|\Gamma_{l}|>k_{l}$ and $|\Gamma_{3-l}|\leq k_{3-l}$ for $l\in\{1,2\}$. Then, our algorithm simply proceeds as follows:
  • Construct $\Gamma_{l}^{\prime}=\{i|i\in\Gamma_{l},d\left(i,\Gamma_{3-l}\right)>3r^{*}\}$ and
    return $C=\Gamma_{3-l}\cup\Gamma_{l}^{\prime}$ as the desired center set.
The correctness of such $C$ can be derived from the following theorem: